OpenStudy (anonymous):

Urgent please help as soon as possible In the reaction of 1.18g of salicylic acid (MM= 138.12g/mol) with 2.88g of acetic anhydride (MM= 102.10g/mol) a student obtained 1.08g of acetylsalicyclic acid (MM= 180.17 g/mol). What is the percent yield?

OpenStudy (anonymous):

First, you have to write down the equation for the reaction as it will show you the ratio of the moles from each compound. Then you calculate the moles of each compound first, based on mass and molecular mass. After you find out how much the theoretical mass of the product is, you divide the mass obtained from the experiment by the theoretical mass.

OpenStudy (anonymous):

Can't write the equation here but I assume you can manage with it, yourself. What it shows is that the molar ratio of acid:anhydride:product = 1:1:1. Therefore, an equal number of moles can react to produce the desired compound. Here are the calculations, themselves: n(acid) = 1.18 / 138.12 = 0.0085 mol n(anhydride) = 2.88 / 102.10 = 0.028 mol What we notice is that the anhydride is in excess. This is done due to practical reasons as we would like as much of the other reagent to react as possible. So, the moles of the product that we're theoretically supposed to get are 0.0085 mol. We have the molecular mass, so the theoretical mass is m(ThProd) = 0.0085 * 180,17 = 1.531 g. In practice the student got m(PrProd) = 1.08 g Which means this is 1.08 / 1.153 * 100 = 93.67 %. Pretty good job he did.